# Are AI-Generated Fixes Secure? Analyzing LLM and Agent Patches on SWE-bench

### v2

@@ III-C1 Code-Level Factors
Files Added or Modified (All) 1.72 1.25 $<$0.001 0.225
Files Added or Modified (Python) 1.60 1.14 $<$0.001 0.234
Unique File Types Modified 1.11 1.00 0.0002 0.096
Sensitive Files Modified 0.005 0.007 0.919 -0.001
LOC Generated by LLM (Python) 106.56 81.84 $<$0.001 0.265
Avg. Cyclomatic Complexity of LLM-Generated Code 2.88 3.03 0.7051 -0.019
TABLE III: Summary statistics comparing vulnerable vs. all Llama/agentic-generated code for code-level metrics.
@@ V A V A V A
Files Added or Modified (All) 158.40 11.38 1.33 1.07 1.67 1.46
Files Added or Modified (Python) 68.83 4.46 1.33 1.07 1.67 1.40
Unique File Types Modified 2.79 1.41 1.00 0.98 1.00 0.90
Sensitive Files Modified 0.191 0.013 0.00 0.00 0.00 0.004
LOC Generated by LLM (Python) 997.98 1614.53 387.33 1294.69 2406.00 1764.98
Avg. Cyclomatic Complexity of LLM-Generated Code 2.79 1.41 1.00 0.98 1.00 0.90
TABLE IV: Mean values for vulnerable (V) vs. all (A) agentic-generated patches across code-level metrics.
# Files Added or Modified: As shown in Table III, Llama-generated vulnerable patches (Vuln Mean) involve modifications to a larger number of files compared to the entire LLM-generated set (All Mean), both in overall (1.72 vs. 1.25) and within Python files (1.60 vs. 1.14), with statistically significant differences (p $<$ 0.001, Cliff’s $\delta$ = 0.23).
Among the agentic frameworks, OH showed an even more dramatic pattern: vulnerable instances modified 158.40 files on average (vs. 11.38), including Python files (68.83 vs. 4.46), with effect size of $\delta$=0.42 and a p-value$<$0.001. This reflects the framework’s tendency to perform excessive rewrites. Additionally, the differences in ACR (1.33 vs. 1.07) and HC (1.67 vs. 1.46) showed higher means for vulnerable instances.
@@
# Unique File Types Modified: Vulnerable code modifications by Llama affected a slightly broader range of file types (mean = 1.11 vs. 1.00). This difference is statistically significant (p = 0.0002), although the effect size is negligible (Cliff’s $\delta$ = 0.096). This suggests that file type diversity alone is a weak indicator of vulnerability risk but may still inform broader risk assessments when combined with other features.
@@
Across the frameworks, we observe similar patterns: ACR (1.00 vs. 0.98) and HC (1.00 vs. 0.90) both show higher mean values for the vulnerable instances. Further, OH exhibited a more notable increase (2.79 vs. 1.41), with statistical significance and meaningful effect size ($\delta$ = 0.153, p = 0.0031). This reinforces earlier findings about OH’s broader file editing behavior, which may contribute to higher vulnerability risk in its generated patches.
@@
Figure 2: Number of changed files per instance (capped at 10). Vulnerable cases show greater variability and significantly higher averages than the general set.
@@
# Sensitive file modifications: using our predefined a set of sensitive file types (e.g., .conf, .yml, .json), we compare their modification rates in vulnerable and all instances. In Llama-generated code vulnerable instances did not show significant changes in modifying such files (p = 0.919), indicating that LLMs do not disproportionately touch sensitive configuration files when generating insecure code.
@@
This pattern holds for ACR (0.0 vs. 0.0) and HC (0.0 vs. 0.004). OH, on the other hand, exhibited a statistically significant increase in sensitive file edits (0.19 vs. 0.01, p $<$ 0.001), suggesting that its more expansive editing behavior may lead it to touch more sensitive files. However, the effect size remained negligible ($\delta$ = 0.076), and results across frameworks were mixed. This indicates that while sensitive file edits are a contributing factor in certain frameworks, they are not a consistent predictor of vulnerability on their own.
LOC Generated by LLM: This factor represents the scale of code generation by the LLM. In the Llama-generated patches, vulnerable instances involve more extensive edits (106.6 vs. 81.8 lines on average), a statistically significant difference (p $<$ 0.001) with a positive effect size ($\delta$ = 0.265). This suggests that larger code generations are more prone to vulnerabilities.
@@
Among the agentic frameworks, ACR shows the strongest contrast, with vulnerable outputs averaging fewer lines (387 vs. 1295). OH also shows a reduction in generated LOC for vulnerable instances (998 vs. 1615), with a small effect size ($\delta$ = -0.148, p = 0.082). In contrast, HC exhibits the opposite trend, with vulnerable code averaging more lines (2406 vs. 1765). These contrasting patterns suggest that while the scale of generation is associated with vulnerability likelihood, the relationship can be highly dependent on the specific framework’s behavior and other contextual factors.
@@
Cyclomatic Complexity of LLM-Generated Code: We assess whether the vulnerable outputs contain more complex control flows by measuring the average cyclomatic complexity of functions in files generated or modified by the LLM. In the Llama patches, vulnerable instances show slightly lower complexity (2.88 vs. 3.03), but this difference is not statistically significant (p = 0.7051, $\delta$ = -0.019), suggesting that structurally simple code may still contain security flaws.
@@
Among agentic frameworks, OH shows a similar trend, with vulnerable files having lower complexity (3.63 vs. 4.51), though the difference is not statistically significant (p = 0.074, $\delta$ = -0.085). Further, HC shows a slight increase in the complexity of vulnerable files (5.28 vs. 4.36). Notably, ACR exhibits a significant increase in complexity of vulnerable files (7.35 vs. 4.30). Our observations highlight that vulnerability is not limited to structurally complex code as simple logic can still be insecure. Yet, abnormally high complexity, as with the patches of ACR, can serve as an indicator for security risk, offering a practical signal for prioritizing code review.
@@
Code-Level Factors Across all systems, vulnerable patches typically modify more files (Llama, OH) and file types (OH), and involve more generated code (Llama, HC). Higher-risk patches in ACR exhibit abnormally high code complexity.
III-C2 Issue-Level Factors
@@
Issue Type and Bug Type Classification: Among all the vulnerable instances generated by LLama and the agentic frameworks, the vast majority of issues were labeled as Bug. For LLama-generated vulnerable patches, 138 out of 185 (73.4%) were classified as Bug, 40 (21.27%) as Feature, and 3 (1.59%) as Question. In contrast, OH had 10 Bugs and 5 Feature; ACR had 2 Bugs and 1 Feature; and HC had 2 Bugs.
Table V summarizes the distribution of bug types across all vulnerable instances. The most frequent categories are Program Anomaly and Configuration. Other bug types, such as Performance, GUI-related, appear less frequently. Overall, vulnerabilities of the standalone LLM and agentic frameworks mostly involve addressing core logic or environment-related issues, where deeper contextual understanding is critical.
@@
Program Anomaly Issue 82 7 2 1
Configuration Issue 27 2
Performance Issue 6
GUI-Related Issue 4
Test Code-Related Issue 3 1
@@
Security Issue 2 1
Network Issue 1
Miscellaneous/Other Issue 11
TABLE V: Bug type distribution across vulnerable instances.
@@
Information Completeness: Among vulnerable LLM cases, nearly all (136/138) bug-related issues included observed behavior, but just about 65.9% (91/138) included steps to reproduce, and only 49.2% (68/138) provided expected behavior. These rates indicate that while problems are usually described, some important information, i.e., what the code is supposed to do and how to reproduce the issue, is often missing. This lack of information may increase ambiguity, forcing the model to infer expected behavior without explicit instructions, thereby raising the risk of inadequate implementations.
@@
Examining the 18 unique GitHub issues for agentic workflows (15 for OH, 3 for ACR, 2 for HC), we observe a high rate of information completeness for OH instances: all 10 vulnerable issue reports included observed behavior, 8 included expected behavior, and 9 provided steps to reproduce. In ACR, both vulnerable instances (2/2) had all three types of information. For HC, 2 out of 3 issues include observed behavior, 1 provides expected behavior, and 2 offer reproduction steps. While sample sizes are limited, we observe that framework-generated vulnerabilities can occur even when issue descriptions are reasonably specified.
@@
# Words: We observed minor differences in the length of issue descriptions and associated comments in vulnerable vs. all Llama-generated instances. On average, problem descriptions contained 111.71 words for all issues and 112.55 words for vulnerable instances. Neither the standalone word count of problem descriptions (p = 0.753, Cliff’s $\delta$ = 0.013) nor the combined word count including comments (p = 0.191, $\delta$ = 0.056) reached statistical significance. These results suggest that the overall length of issue content alone is not a strong differentiator in predicting whether the model will introduce vulnerabilities, and possibly other factors, such as clarity or specificity, can play a more important role than sheer verbosity.
@@
Across the frameworks, we observed mixed patterns. OH and ACR had higher average word counts when combining issue and comment text (355.1 and 435.7 words, respectively), while HC’s was much lower (96.0). This illustrates the unique way each framework is impacted by the available information.
@@
# Comments Prior to First PR: Llama-generated vulnerable instances had slightly more comments on average (4.14) compared to all issues (3.57), but the difference was not significant (p = 0.22, $\delta$ = 0.053), indicating that comment count does not possibly distinguish vulnerable from non-vulnerable LLM-generated patches.
Among the frameworks, OH had significantly fewer comments than the general set (p $<$ 0.001, $\delta$ = -0.315), while ACR and HC showed no statistically significant differences (p = 0.6256 and p = 0.2324, respectively). While the number of comments is not a strong predictor overall, certain frameworks may be more sensitive to the availability, or absence, of prior discussion when generating code.
@@
Presence of Code Snippets: We observe that only 48.1% of issues associated with vulnerable LLM outputs include code snippets in the issue body, compared to 58.6% of all issues. This difference is statistically significant ($\chi^{2}$ = 7.85, p = 0.0051), suggesting that the presence or absence of code snippets is not evenly distributed between the two groups. The chi-squared test evaluates whether there is a meaningful association between two categorical variables, in this case, code snippet presence and vulnerability occurrence. To assess the strength of this association, we calculate the phi coefficient ($\phi$ = 0.019), a standardized effect size for 2×2 contingency tables. While statistically significant, the effect size is negligible, indicating a weak association. Nonetheless, the direction of the effect suggests that LLMs are more prone to generating vulnerabilities when issue descriptions lack code snippets, possibly due to the absence of concrete examples, stack traces, or code context, which increases ambiguity.
OH shows an even stronger pattern ($\chi^{2}$ = 32.29, p $<$ 0.0001, $\phi$ = 0.039), suggesting that missing code context may be particularly harmful for more autonomous systems. While ACR had no vulnerable issues with code snippets (0 out of 3), both HC issues contained code snippets (2 out of 2). Overall, these trends indicate that the potential ambiguities in issues lacking code snippets can increase the risk of vulnerabilities.
@@
Issue-Level Factors Insecure patches are associated with issues where code snippets are absent, and when expected behavior or reproduction steps are missing. Most of these issues are labeled as Bug and involve core logic or configuration problems. OH is particularly sensitive to issues without code snippets.
III-C3 Project-Level Factors
@@
# Files in Repository: We examined whether the size of the codebase, as measured by the total number of Python files in the repository, correlates with vulnerability outcomes. For Llama-generated code, the correlation was weak and not statistically significant ($\rho=0.188$, p = 0.206). Similarly, we found no meaningful correlations in any of the agentic frameworks: OH ($\rho=0.171$, p = 0.251), ACR ($\rho=0.205$, p = 0.168), and HC ($\rho=0.171$, p = 0.251). These findings suggest that codebase size is not a reliable indicator of LLM-generated vulnerabilities.
@@
Maintainability and Complexity: Similarly, we found no statistically significant correlation between vulnerability counts and average maintainability index or the cyclomatic complexity per function in the projects. Within the LLM-generated code, maintainability ($\rho=-0.082$, p = 0.583) and complexity ($\rho=-0.064$, p = 0.670) were not correlated with vulnerability introduction. The same holds for the frameworks. For example, in OH, maintainability ($\rho=0.098$, p = 0.512) and complexity ($\rho=-0.007$, p = 0.962) were not predictive of vulnerability counts. ACR and HC showed similar patterns. These findings indicate that the internal complexity or maintainability of a project is not a strong predictor of insecure code generation by LLM-based systems.
@@
Contributor Count: We also examined whether the number of contributors in a project correlates with the introduction of vulnerabilities. For LLM-generated code, the association was weak and statistically insignificant ($\rho=0.166$, p = 0.2635). The same held true across all agentic frameworks: OH ($\rho=0.161$, p = 0.280), ACR ($\rho=0.158$, p = 0.288), and HC ($\rho=0.109$, p = 0.467). These findings suggest that community size, used here as a rough proxy for project activity, does not significantly influence LLM vulnerability outcomes, although the consistently positive (albeit weak) correlations hint at a very slight directional trend.
@@
Project-Level Factors Project size, complexity, and contributor count showed weak, non-significant correlations with vulnerability rates. As a result, these factors alone do not reliably predict insecure LLM-generated code.
RQ3: Summary of Findings. Our analysis of over 20,000 LLM-generated patches reveals that vulnerability-prone outputs are associated with a consistent set of characteristics, particularly at the code and issue level. At the code level, vulnerable patches tend to involve broader and more scattered edits, modifying more files and generating more lines of code, especially in more autonomous frameworks like OH. While unique file types and sensitive file modifications showed some variation, their predictive value remains limited across all systems. Cyclomatic complexity was not a strong differentiator of insecure code, except in ACR, where vulnerable patches had significantly higher complexity.
At the issue level, vulnerabilities were frequently found in tasks labeled as bugs, particularly those lacking full contextual descriptions (e.g., missing expected behavior or reproduction steps). However, for both the standalone LLM and frameworks, higher word counts did not strongly distinguish vulnerable from non-vulnerable cases. A key trend was that issues lacking concrete code snippets were more frequently linked to insecure patches, highlighting the importance of prompts with specific code snippets for secure patch generation.
At the project level, no strong correlation was found between vulnerability introduction and repository size, complexity, maintainability, or contributor count. These results suggest that security risks are impacted more by how LLMs interact with individual tasks and contextual cues than by the broader structural properties and attributes of the repository.
Overall, vulnerabilities in LLM-generated patches are most strongly associated with code-level factors like broader edits and file changes, and issue-level factors such as missing contextual information, especially the absence of code snippets. In contrast, project-level attributes show no clear relationship with vulnerability outcomes.
IV Threats to Validity

### v3

@@ 3.3.1. Code-Level Factors
Files Added or Modified (All) 1.72 1.25 <<0.001 0.200
Files Added or Modified (Python) 1.60 1.16 <<0.001 0.199
Unique File Types Modified 1.10 0.99 0.007 0.081
Sensitive Files Modified 0.00 0.007 0.369 -0.006
LOC Generated by LLM (Python) 1740.50 2218.12 <<0.001 0.197
Avg. Cyclomatic Complexity of LLM-Generated Code 2.88 3.02 0.8796 -0.009
Table 3. Summary statistics comparing vulnerable vs. all Llama/agentic-generated code for code-level metrics
@@ V A V A V A
Files Added or Modified (All) 344.27 11.38 1.33 1.07 1.50 1.46
Files Added or Modified (Python) 68.83 4.46 1.33 1.07 1.50 1.40
Unique File Types Modified 3.73 1.41 1.00 0.98 1.00 0.90
Sensitive Files Modified 0.0909 0.0136 0.00 0.00 0.00 0.004
LOC Generated by LLM (Python) 1218.6 1591.0 387.33 1294.6 1591 1764
Avg. Cyclomatic Complexity of LLM-Generated Code 4.015 4.513 7.352 4.303 6.470 4.363
Table 4. Mean values for vulnerable (V) vs. all (A) agentic-generated patches across code-level metrics.
# Files Added or Modified: As shown in Table 3, Llama-generated vulnerable patches (Vuln Mean) involve modifications to a larger number of files compared to the entire LLM-generated set (All Mean), both in overall (1.72 vs. 1.25) and within Python files (1.60 vs. 1.16), with statistically significant differences (p << 0.001, Cliff’s $\delta$ = 0.2 & 0.19).
Among the agentic frameworks, OH showed an even more dramatic pattern: vulnerable instances modified 344.27 files on average (vs. 11.38), including Python files (68.83 vs. 4.46), with large effect size of $\delta$=0.77 and a p-value<<0.001. This reflects the framework’s tendency to perform excessive rewrites. Additionally, the differences in ACR (1.33 vs. 1.07) and HC (1.50 vs. 1.46) showed higher means for vulnerable instances.
@@
# Unique File Types Modified: Vulnerable code modifications by Llama affected a slightly broader range of file types (mean = 1.10 vs. 0.99). This difference is statistically significant (p = 0.007), although the effect size is negligible (Cliff’s $\delta$ = 0.096). This suggests that file type diversity alone is a weak indicator of vulnerability risk but may still inform broader risk assessments when combined with other features.
@@
Across the frameworks, we observe similar patterns: ACR (1.00 vs. 0.98) and HC (1.00 vs. 0.90) both show higher mean values for the vulnerable instances. Further, OH exhibited a more notable increase (3.73 vs. 1.41), with statistical significance and meaningful effect size ($\delta$ = 0.345, p = 0.0011). This reinforces earlier findings about OH’s broader file editing behavior, which may contribute to higher vulnerability risk in its generated patches.
@@
Figure 2. Number of changed files per instance (capped at 10). Vulnerable cases show greater variability and significantly higher averages than the general set.
@@
# Sensitive file modifications: using our predefined a set of sensitive file types (e.g., .conf, .yml, .json), we compare their modification rates in vulnerable and all instances. In Llama-generated code vulnerable instances did not show significant changes in modifying such files (p = 0.369), indicating that Llama 3.3 does not disproportionately touch sensitive configuration files when generating insecure code.
@@
This pattern holds for ACR (0.0 vs. 0.0) and HC (0.0 vs. 0.004). OH, on the other hand, exhibited a statistically significant increase in sensitive file edits (0.09 vs. 0.01, p << 0.008), suggesting that its more expansive editing behavior may lead it to touch more sensitive files. However, the effect size remained negligible ($\delta$ = 0.081), and results across frameworks were mixed. This indicates that while sensitive file edits are a contributing factor in certain frameworks, they are not a consistent predictor of vulnerability on their own.
LOC Generated by LLM: This factor represents the scale of code generation by the LLM. In the Llama-generated patches, vulnerable instances involve less extensive edits (1740 vs. 2218 lines on average), a statistically significant difference (p << 0.001) with a positive effect size ($\delta$ = 0.197). This suggests that larger code generations are not more prone to vulnerabilities. As demonstrated in prior work, even highly concise edits can introduce substantial security risks. For instance, severe vulnerabilities such as Heartbleed and ShellShock stemmed from only a handful of modified lines, showing how small changes to core logic can have extreme consequences (lovgren2023can). This findings also aligns with the empirical studies of developer activity, revealing that vulnerability-inducing commits are shrinking in size over time, with smaller commits associated with more severe vulnerabilities (andrade2021investigating; jiang2024understanding). Viewed in this context, our finding suggests that the brevity of LLM-generated patches may reflect the models’ difficulty in fully reasoning about program context. When models alter a few lines of critical logic based on incomplete assumptions, these compact changes can still meaningfully compromise the security of software.
@@
Among the agentic frameworks, ACR shows the strongest contrast, with vulnerable outputs averaging fewer lines (387 vs. 1294). OH also shows a reduction in generated LOC for vulnerable instances (1218 vs. 1591), with a small effect size ($\delta$ = -0.10, p = 0.51). In contrast, HC exhibits the opposite trend, with vulnerable code averaging more lines (1764 vs. 1591). These contrasting patterns suggest that while the scale of generation can be associated with vulnerability likelihood, the relationship can be highly dependent on the specific framework’s behavior and other contextual factors.
@@
Cyclomatic Complexity of LLM-Generated Code: We assess whether the vulnerable outputs contain more complex control flows by measuring the average cyclomatic complexity of functions in files generated or modified by the LLM. In the Llama patches, vulnerable instances show slightly lower complexity (2.88 vs. 3.02), but this difference is not statistically significant (p = 0.8796, $\delta$ = -0.009), suggesting that structurally simple code may still contain security flaws.
@@
Among agentic frameworks, OH shows a similar trend, with vulnerable files having lower complexity (4.015 vs. 4.513), though the difference is not statistically significant (p = 0.688, $\delta$ = 0.035). Further, HC shows a slight increase in the complexity of vulnerable files (6.47 vs. 4.363). Notably, ACR exhibits a significant increase in complexity of vulnerable files (7.352 vs. 4.303). Our observations highlight that vulnerability is not limited to structurally complex code as simple logic can still be insecure. Yet, abnormally high complexity, as with the patches of ACR, can serve as an indicator for security risk, offering a practical signal for prioritizing code review.
@@
Code-Level Factors Across all systems, vulnerable patches typically modify more files and file types, but involve fewer lines of generated code. Higher-risk patches in ACR exhibit abnormally high code complexity.
3.3.2. Issue-Level Factors
@@
Issue Type and Bug Type Classification: Among all the vulnerable instances generated by Llama and the agentic frameworks, the vast majority of issues were labeled as Bug. For Llama-generated vulnerable patches, 105 out of 135 (77.7%) were classified as Bug, 29 (21.4%) as Feature, and 1 (0.7%) as Question. In contrast, OH had 9 Bugs and 35 Feature; ACR had 2 Bugs and 1 Feature; and HC had 2 Bugs. Here, OH stands out as it often generated vulnerable code when attempting to resolve issues that seek to implement new feature within projects.
Table 5 summarizes the distribution of bug types across all vulnerable instances. The most frequent categories are Program Anomaly and Configuration. Other bug types, such as Performance, GUI-related, appear less frequently. Overall, vulnerabilities of Llama 3.3 and the agentic frameworks mostly involve addressing core logic or environment-related issues, where deeper contextual understanding is critical.
@@
Program Anomaly Issue 63 7 1 2
Configuration Issue 18 2
Miscellaneous/Other Issue 8
Performance Issue 5
GUI-Related Issue 3
Test Code-Related Issue 3
Network Issue 3
@@
Security Issue 1
Table 5. Bug type distribution across vulnerable instances.
@@
Information Completeness: Among vulnerable LLM cases, nearly all (101/105) bug-related issues included observed behavior, but just about 64.7% (68/105) included steps to reproduce, and only 39% (41/105) provided expected behavior. These rates indicate that while problems are usually described, some important information, i.e., what the code is supposed to do and how to reproduce the issue, is often missing. This lack of information may increase ambiguity, forcing the model to infer expected behavior without explicit instructions, thereby raising the risk of inadequate implementations.
@@
Examining the 11 unique GitHub issues for agentic workflows (6 for OH, 3 for ACR, 2 for HC), we observe a high rate of information completeness for OH instances: all vulnerable issue reports included observed and expected behavior and provided steps to reproduce. In ACR, both vulnerable instances (2/2) had all three types of information. For HC, 2 out of 3 issues include observed behavior, 1 provides expected behavior, and 2 offer reproduction steps. While sample sizes are limited, we observe that framework-generated vulnerabilities can occur even when issue descriptions are reasonably specified.
@@
# Words: We observed minor differences in the length of issue descriptions and associated comments in vulnerable vs. all Llama-generated instances. On average, problem descriptions contained 111.71 words for all issues and 105.78 words for vulnerable instances. Neither the word count of problem descriptions alone (p = 0.5346, Cliff’s $\delta$ = -0.031) nor the combined word count including comments (p = 0.4726, $\delta$ = 0.036) reached statistical significance. These results suggest that the overall length of issue content alone is not a strong differentiator in predicting whether the model will introduce vulnerabilities, and possibly other factors, such as clarity or specificity, can play a more important role than the amount of available textual data.
@@
Across the frameworks, we observed mixed patterns. OH and ACR had higher average word counts when combining issue and comment text (547.16 and 435.67 words, respectively), while HC’s was much lower (96.0). This illustrates the unique way each framework is impacted by the available information.
@@
# Comments Prior to First PR: Llama-generated vulnerable instances had slightly more comments on average (4.1) compared to all issues (3.6), but the difference was not significant (p = 0.68, $\delta$ = 0.025), indicating that comment count does not clearly distinguish vulnerable from non-vulnerable LLM-generated patches.
Across all three frameworks, none of the vulnerable issues contained comments prior to the first pull request. This absence suggests that, unlike in the Llama-generated set, prior discussion was not associated with an increase in the vulnerabilities observed within the frameworks.
@@
Presence of Code Snippets: We observe that only 47.4% of issues associated with vulnerable LLM outputs include code snippets in the issue body, compared to 58.6% of all issues. This difference is statistically significant ($\chi^{2}$ = 6.44, p = 0.011), suggesting that the presence or absence of code snippets is not evenly distributed between the two groups. The chi-squared test evaluates whether there is a meaningful association between two categorical variables, in this case, code snippet presence and vulnerability occurrence. To assess the strength of this association, we calculate the phi coefficient ($\phi$ = 0.017), a standardized effect size for 2×2 contingency tables. While statistically significant, the effect size is negligible, indicating a weak association. Nonetheless, the direction of the effect suggests that Llama 3.3 is more prone to generating vulnerabilities when issue descriptions lack code snippets, possibly due to the absence of concrete examples, stack traces, or code context, which increases ambiguity. Overall, these trends indicate that the potential ambiguities in issues that lack code snippets can increase the risk of vulnerabilities.
@@
Issue-Level Factors Insecure patches from Llama 3.3 are associated with issues where code snippets are absent, and when expected behavior or reproduction steps are missing. Most of these issues are labeled as Bug and involve core logic or configuration problems. OH is particularly sensitive to issues related to feature development and those without code snippets.
3.3.3. Project-Level Factors
@@
# Files in Repository: We examined whether the size of the codebase, as measured by the total number of Python files in the repository, correlates with vulnerability outcomes. For Llama-generated code, the correlation was weak and not statistically significant ($\rho=0.136$, p = 0.362). Similarly, we found no meaningful correlations in any of the agentic frameworks: OH ($\rho=0.70$, p = 0.642), ACR ($\rho=0.205$, p = 0.168), and HC ($\rho=0.171$, p = 0.251). These findings suggest that codebase size is not a reliable indicator of LLM-generated vulnerabilities.
@@
Maintainability and Complexity: Similarly, we found no statistically significant correlation between vulnerability counts and average maintainability index or the cyclomatic complexity per function in the projects. Within the LLM-generated code, maintainability ($\rho=-0.006$, p = 0.968) and complexity ($\rho=-0.062$, p = 0.679) were not correlated with vulnerability introduction. The same holds for the frameworks. For example, in OH, maintainability ($\rho=-0.021$, p = 0.888) and complexity ($\rho=0.049$, p = 0.744) were not predictive of vulnerability counts. ACR and HC showed similar patterns. These findings indicate that the internal complexity or maintainability of a project is not a strong predictor of insecure code generation by LLM-based systems.
@@
Contributor Count: We also examined whether the number of contributors in a project correlates with the introduction of vulnerabilities. For Llama-generated code, the association was weak and statistically insignificant ($\rho=0.184$, p = 0.217). The same held true across all agentic frameworks: OH ($\rho=0.210$, p = 0.156), ACR ($\rho=0.158$, p = 0.288), and HC ($\rho=0.109$, p = 0.467). These findings suggest that community size, used here as a rough proxy for project activity, does not significantly influence LLM vulnerability outcomes, although the consistently positive (albeit weak) correlations hint at a very slight directional trend.
@@
Project-Level Factors Project size, complexity, and contributor count showed weak, non-significant correlations with vulnerability rates. As a result, these factors alone were not observed to reliably predict insecure LLM-generated code.
RQ3 Summary: Across all systems, vulnerable patches are most strongly associated with code and issue-level characteristics rather than project-level properties. At the code level, insecure patches consistently modify more files (and, for OH, more file types), while often involving relatively small, localized edits in each file. In ACR, vulnerable patches also exhibit substantially higher cyclomatic complexity than non-vulnerable ones. At the issue level, vulnerabilities predominantly present in bug reports and configuration-related issues, especially when expected behavior and reproduction steps are under-specified and when issue descriptions lack concrete code snippets. In contrast, project-level factors such as repository size, cyclomatic complexity, and contributor count show non-significant correlations with vulnerability counts.
4. Threats to Validity